# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ C.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
Proof of Theorem 2 : Without loss of generation, we assume groups $1,2,3$ of $\Lambda^{*}$ satisfy Condition 5 and $|\mathcal{B}_{5}^{*}|=|\mathcal{B}_{6}^{*}|=|\mathcal{B}_{7}^{*}|=2$. The proof consists of three part: (1) show $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$;(2) show there exists some $D\in\mathcal{D}$ and $P\in\mathcal{P}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]PD$ and $(PD\Phi DP^{\top})[\{1,\ldots,5\},\{1,\ldots,5\}]=\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]$. (3) show $\Lambda[\cup_{i=5}^{7}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=5}^{7}\mathcal{B}^{*}_{i},:]PD$, $(PD\Phi DP^{\top})[\{6,7,8\},\{6,7,8\}]=\Phi^{*}[\{6,7,8\},\{6,7,8\}]$ and $\psi_{i}=\psi_{i}^{*}$ for $i\in\mathcal{B}_{5},\ldots,\mathcal{B}_{7}$.
We first show that $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$. According to the proof of Theorem 5.1 of Anderson and Rubin, (1956), for any $j\in\{1,\ldots,J\}$, if the $j$th row $\Lambda^{*}$ is deleted and there exist two disjoint submatrices of the rest of $\Lambda^{*}$ with full columns rank, then $\psi_{j}=\psi_{j}^{*}$. We now construct the partitions for $j\in\cup_{i=1}^{4}\mathcal{B}^{*}_{i}$. By Condition 5, there exist 3 different observations $i_{k1},i_{k2},i_{k3}\in\mathcal{B}_{k}^{*}$ such that $\Lambda^{*}[\{i_{k1}\},\{1+k\}]$,$\Lambda^{*}[\{i_{k2}\},\{1+k\}]$ and $\Lambda^{*}[\{i_{k3}\},\{1+k\}]\neq 0$ for $k=1,2,3,4$. Furthermore, there exist 2 different observations $i_{k4},i_{k5}\in\mathcal{B}_{k}^{*}$ such that $\Lambda^{*}[\{i_{k4},i_{k5}\},\{1,1+k\}]$ has rank 2 for $k=1,2,3$. By Condition 4, there exist 2 different observations $\Lambda^{*}[\{i_{k2}\},\{1+k\}]$ and $\Lambda^{*}[\{i_{k2}\},\{1+k\}]\neq 0$ for $k=5,6,7$.
Now for $j\in\mathcal{B}^{*}_{4}$, for example $j=i_{4,3}$, we select the first submatrix that consists of rows $i_{14},i_{15},i_{21},i_{31},\ldots,\\ i_{71}$ of $\Lambda^{*}$ and the second submatrix that consists of rows $i_{11},i_{24},i_{25},i_{32},\ldots,i_{72}$, where $i_{11}\neq i_{14},i_{15}$ and $i_{21}\neq i_{24},i_{25}$. It is easy to check that the two submatrices have column rank 8. For $j\in\mathcal{B}_{1}^{*}$, for example $j=i_{11}$, we construct the first submatrix that consists of rows $i_{12},i_{24},i_{25},i_{31},i_{41},\ldots,i_{71}$ of $\Lambda^{*}$, where $i_{31}\neq i_{34},i_{35}$ and the second submatrix that consists of rows $i_{13},i_{21},i_{34},i_{35},i_{42},\ldots,i_{72}$ of $\Lambda^{*}$, where $i_{21}\neq i_{24},i_{25}$. It is easy to check that the two submatrices have column rank 8. The construction of submatrices for $j\in\mathcal{B}^{*}_{2},\mathcal{B}^{*}_{3}$ is similar to the case of $j\in\mathcal{B}^{*}_{1}$. Then, by Theorem 5.1 of Anderson and Rubin, (1956), $\psi_{i}=\psi^{*}_{i}$ for $i\in\mathcal{B}_{1},\ldots,\mathcal{B}_{4}$.
@@
Secondly, with a similar argument in Theorem 1, there exists some $D\in\mathcal{D}$ and $P\in\mathcal{P}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]PD$ and $(PD\Phi DP^{\top})[\{1,\ldots,5\},\{1,\ldots,5\}]=\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]$. Without loss of generation, we assume $P$ and $D$ are identity matrices.

### v2

@@ G.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
Proof of Theorem 2 : Without loss of generation, we assume that $|\mathcal{B}_{1}^{*}|=|\mathcal{B}_{2}^{*}|=5$, $|\mathcal{B}_{3}^{*}|=|\mathcal{B}_{4}^{*}|=4$ and $|\mathcal{B}_{5}^{*}|=|\mathcal{B}_{6}^{*}|=|\mathcal{B}_{7}^{*}|=2$. Suppose that there exists $\Lambda$, $\Phi$ and $\Psi$ and $\Sigma=\Lambda\Phi\Lambda^{\top}+\Psi$ such that $\Sigma=\Sigma^{*}$. The proof consists of two parts: (1) Show that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]$ and $\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]$ has the same bi-factor structure. Without loss of generality, we further assume that $\Lambda[\mathcal{B}_{i}^{*},\{1+i\}]\neq\mathbf{0}$ for $i=1,\ldots,4$. We show that there exists some $5\times 5$ sign flip matrix $\widetilde{D}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},\{1,\ldots,5\}]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},\{1,\ldots,5\}]\widetilde{D}$, $\Phi[\{1,\ldots,5\},\{1,\ldots,5\}]=\widetilde{D}\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]\widetilde{D}$ and $\psi_{j}=\psi^{*}_{j}$ for $j\in\cup_{i=1}^{4}\mathcal{B}^{*}_{i}$. (2) Show that $\Lambda$ and $\Lambda^{*}$ have the same bi-factor structure for the rest of the variables and complete the proof.
We now prove the first part. Let $\mathcal{F}_{i}=\{g:\Lambda[\mathcal{B}_{i}^{*},\{1+g\}]\neq\mathbf{0}\}\cup\{1\}$ be the set of factors such that the variables belonging to $\mathcal{B}_{i}^{*}$ load on these factors for $i=1,\ldots,4$. We note that $|\mathcal{F}_{i}|\geq 2$ for $i=1,\ldots,4$. When $|\mathcal{F}_{i}|=2$, all variables that belong to $\mathcal{B}_{i}^{*}$ load on the same group factor. We claim that
[…]
If $|\mathcal{F}_{i}|\leq|\mathcal{B}^{*}_{i}|$, there exists some $g_{i}\in\mathcal{F}_{i}$, $g_{i}\neq 1$ and $j_{g_{i}},j^{\prime}_{g_{i}}\in\mathcal{B}_{i}^{*}$ such that $\lambda_{j_{g_{i}},1+g_{i}}\neq 0$ and $\lambda_{j^{\prime}_{g_{i}},1+g_{i}}\neq 0$. For $1\leq i^{\prime}\leq 4$, $i^{\prime}\neq i$, consider the equation $\Sigma[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\mathcal{B}_{i^{\prime}}^{*}]=\Sigma^{*}[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\mathcal{B}_{i^{\prime}}^{*}]$, which is equivalent to
$\Lambda[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\{1,1+g_{i}\}]\Phi[\{1,1+g_{i}\},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top}$ (G.10)
$=$ $\Lambda^{*}[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\{1,1+i\}]\Phi^{*}[\{1,1+i\},\{1,1+i^{\prime}\}](\Lambda^{*}[\mathcal{B}_{i^{\prime}}^{*},\{1,1+i^{\prime}\}])^{\top}.$
Noticing that by Condition 5 and 6 hold for $\Phi^{*}$ and $\Lambda^{*}$,
$\Lambda^{*}[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\{1,1+i\}]\Phi^{*}[\{1,1+i\},\{1,1+i^{\prime}\}](\Lambda^{*}[\mathcal{B}_{i^{\prime}}^{*},\{1,1+i^{\prime}\}])^{\top}$
has rank 2. Thus, $\Lambda[\{j_{g_{i}},j^{\prime}_{g_{i}}\},\{1,1+g_{i}\}]$ should have rank 2. Otherwise, $\Lambda[\{j_{i},j_{i}^{\prime}\},\{1,1+g_{i}\}]\Phi[\{1,1+g_{i}\},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top}$ has at most rank 1, which contradicts (G.10). Then, since for each $g^{\prime}_{i}\in\mathcal{F}_{i}$, $g^{\prime}_{i}\neq 1$, there exists some $j_{g^{\prime}_{i}}$ such that $\lambda_{j_{g^{\prime}_{i}},g^{\prime}_{i}}\neq 0$, it is easy to check that (G.9) holds.
Then, consider the equation $\Sigma[\mathcal{B}_{i}^{*},\mathcal{B}_{i^{\prime}}^{*}]=\Sigma^{*}[\mathcal{B}_{i}^{*},\mathcal{B}_{i^{\prime}}^{*}]$ for $1\leq i\neq i^{\prime}\leq 4$, which is equivalent to
$\Lambda[\mathcal{B}_{i}^{*},\mathcal{F}_{i}]\Phi[\mathcal{F}_{i},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top}=\Lambda^{*}[\mathcal{B}_{i}^{*},\{1,1+i\}]\Phi[\{1,1+i\},\{1,1+i^{\prime}\}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\{1,1+i^{\prime}\}])^{\top}.$ (G.11)
With the same argument, $\Lambda^{*}[\mathcal{B}_{i}^{*},\{1,1+i\}]\Phi[\{1,1+i\},\{1,1+i^{\prime}\}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\{1,1+i^{\prime}\}])^{\top}$ has rank 2. By Sylvester’s rank inequality, we have
$\text{rank}(\Lambda[\mathcal{B}_{i}^{*},\mathcal{F}_{i}]\Phi[\mathcal{F}_{i},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top})$ (G.12)
$\geq$ $\text{rank}((\Lambda[\mathcal{B}_{i}^{*},\mathcal{F}_{i}])+\text{rank}(\Phi[\mathcal{F}_{i},\mathcal{F}_{i^{\prime}}])+\text{rank}(\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])-|\mathcal{F}_{i}|-|\mathcal{F}_{i^{\prime}}|.$
We consider the following case
  1. 1.
    $|\mathcal{F}_{i}|\leq|\mathcal{B}_{i}^{*}|$ for all $1\leq i\leq 4$. In this case, according to claim (G.9) and Condition 5, inequality (G.12) leads to
    $\text{rank}(\Lambda[\mathcal{B}_{i}^{*},\mathcal{F}_{i}]\Phi[\mathcal{F}_{i},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top})\geq\min(|\mathcal{F}_{i}|,|\mathcal{F}_{i^{\prime}}|),$
    for any $1\leq i\neq i^{\prime}\leq 4$. We then have $\min(|\mathcal{F}_{i}|,|\mathcal{F}_{i^{\prime}}|)=2$. By applying this argument to all pairs $(i,i^{\prime})$, $1\leq i<i^{\prime}\leq 4$, we conclude that there exists at most one $\mathcal{F}_{i}$ such that $|\mathcal{F}_{i}|\geq 3$ for $1\leq i\leq 4$.
    If there exists some $i$ such that $|\mathcal{F}_{i}|\geq 3$ for $i=1,\ldots,4$. Without loss of generality, we assume $|\mathcal{F}_{1}|\geq 3$ and $|\mathcal{F}_{i}|=2$ for $i=2,3,4$. We claim that $|\mathcal{F}_{2}\cup\mathcal{F}_{3}|=3$, in other words, the variables belonging to $\mathcal{B}_{2}^{*}$ and $\mathcal{B}_{3}^{*}$ load on different factors. Otherwise, $|\mathcal{F}_{2}\cup\mathcal{F}_{3}|=2$. Consider the equation
    $\Sigma[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}]=\Sigma^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}],$
    which is equivalent to
    $\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3}]\Phi[\mathcal{F}_{2}\cup\mathcal{F}_{3},\mathcal{F}_{2}\cup\mathcal{F}_{3}](\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3}])^{\top}+\Psi[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}]$ (G.13)
    $=$ $\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}]\Phi^{*}[\{1,3,4\},\{1,3,4\}](\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}])^{\top}+\Psi^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}].$
    Since $\Lambda^{*}$ satisfies Condition 6, noticing that $|\mathcal{B}_{2}^{*}|\geq 4$ and $|\mathcal{B}_{3}^{*}|\geq 4$, it is easy to check that the matrix $\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}]$ satisfies the condition for Theorem 5.1 of Anderson and Rubin, (1956), that is, if any row of $\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}]$ is deleted, there still remains two disjoint submatrices of $\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}]$ with rank 3. By applying Theorem 5.1 of Anderson and Rubin, (1956), we have $\Psi[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}]=\Psi^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}]$. Thus, we further have
    $\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3}]\Phi[\mathcal{F}_{2}\cup\mathcal{F}_{3},\mathcal{F}_{2}\cup\mathcal{F}_{3}](\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3}])^{\top}$ (G.14)
    $=$ $\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}]\Phi^{*}[\{1,3,4\},\{1,3,4\}](\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}])^{\top}.$
    If $|\mathcal{F}_{2}\cup\mathcal{F}_{3}|=2$, then the rank of the matrix in the first line of (G.14) is 2, which contradicts the fact that the rank of the matrix in the second line of (G.14) is 3. Thus, $|\mathcal{F}_{2}\cup\mathcal{F}_{3}|=3$. We note that with a similar argument used in (G.13) and (G.14), we also have $|\mathcal{F}_{2}\cup\mathcal{F}_{4}|=3$, $|\mathcal{F}_{3}\cup\mathcal{F}_{4}|=3$ and $|\cup_{i=2,3,4}\mathcal{F}_{i}|=4$. Then, consider the equation
    $\Sigma[\mathcal{B}_{1}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}]=\Sigma^{*}[\mathcal{B}_{1}^{*},\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*}],$
    which is equivalent to
    $\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}]\Phi[\mathcal{F}_{1},\mathcal{F}_{2}\cup\mathcal{F}_{3}](\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3})^{\top}$ (G.15)
    $=$ $\Lambda^{*}[\mathcal{B}_{1}^{*},\{1,2\}]\Phi^{*}[\{1,2\},\{1,3,4\}](\Lambda^{*}[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\{1,3,4\}])^{\top}.$
    We note that the rank of the matrix in the second line of (G.15) is 2. According to Sylvester’s rank inequality
    $\text{rank}(\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}]\Phi[\mathcal{F}_{1},\mathcal{F}_{2}\cup\mathcal{F}_{3}](\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3})^{\top})$ (G.16)
    $\geq$ $\text{rank}(\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}])+\text{rank}(\Phi[\mathcal{F}_{1},\mathcal{F}_{2}\cup\mathcal{F}_{3}])+\text{rank}(\Lambda[\mathcal{B}_{2}^{*}\cup\mathcal{B}_{3}^{*},\mathcal{F}_{2}\cup\mathcal{F}_{3}])-|\mathcal{F}_{1}|-3$
    $=$ $|\mathcal{F}_{1}|+\min(|\mathcal{F}_{1}|,3)+3-|\mathcal{F}_{1}|-3$
    $=$ $3,$
    which contradicts (G.15).
    Thus, in the case, $|\mathcal{F}_{i}|=2$ for $i=1,\ldots,4$. Consider the equation
    $\Sigma[\cup_{i=1,\ldots,4}\mathcal{B}_{i}^{*},\cup_{i=1,\ldots,4}\mathcal{B}_{i}^{*}]=\Sigma^{*}[\cup_{i=1,\ldots,4}\mathcal{B}_{i}^{*},\cup_{i=1,\ldots,4}\mathcal{B}_{i}^{*}].$ (G.17)
@@
    By the similar argument discussed in (G.13), (G.14) and further applying Theorem 1 to (G.17), we conclude that in this case, $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]$ and $\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},:]$ has the same bi-factor structure. Without loss of generality, we further assume that $\mathcal{F}_{i}=\{1,1+i\}$ for $i=1,\ldots,4$. Then, there exists some $5\times 5$ sign flip matrix $\widetilde{D}$ such that $\Lambda[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},\{1,\ldots,5\}]=\Lambda^{*}[\cup_{i=1}^{4}\mathcal{B}^{*}_{i},\{1,\ldots,5\}]\widetilde{D}$, $\Phi[\{1,\ldots,5\},\{1,\ldots,5\}]=\widetilde{D}\Phi^{*}[\{1,\ldots,5\},\{1,\ldots,5\}]\widetilde{D}$ and $\psi_{j}=\psi^{*}_{j}$ for $j\in\cup_{i=1}^{4}\mathcal{B}^{*}_{i}$.
  2. 2.
    There exists some $1\leq i\leq 4$ such that $|\mathcal{F}_{i}|=1+|\mathcal{B}_{i}^{*}|\geq 5$. In this case, according to (G.12)
    $\text{rank}(\Lambda[\mathcal{B}_{i}^{*},\mathcal{F}_{i}]\Phi[\mathcal{F}_{i},\mathcal{F}_{i^{\prime}}](\Lambda[\mathcal{B}_{i^{\prime}}^{*},\mathcal{F}_{i^{\prime}}])^{\top})\geq 3\mbox{~{}~{}if~{}~{}}|\mathcal{F}_{i^{\prime}}|\geq 4.$ (G.18)
    Thus, $|\mathcal{F}_{i^{\prime}}|\leq 3<|\mathcal{B}^{*}_{i^{\prime}}|$ for all $1\leq i^{\prime}\leq 4$, $i^{\prime}\neq i$. Without loss of generality, let $i=1$. For $i^{\prime}=2,3,4$, by the same argument in case 1, we have $\mathcal{F}_{2}=\{1,1+g_{2}\}$, $\mathcal{F}_{3}=\{1,1+g_{3}\}$ and $\mathcal{F}_{4}=\{1,1+g_{4}\}$ for different $g_{2},g_{3}$ and $g_{4}$. Moreover, $\text{rank}(\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\cup_{i=2,3,4}\mathcal{F}_{i}])=4$.
    Then, consider the equation $\Sigma[\mathcal{B}^{*}_{1},\cup_{i=2,3,4}\mathcal{B}_{i}^{*}]=\Sigma^{*}[\mathcal{B}^{*}_{1},\cup_{i=2,3,4}\mathcal{B}_{i}^{*}]$, which is equivalent to
    $\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}]\Phi[\mathcal{F}_{1},\cup_{i=2,3,4}\mathcal{F}_{i}](\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\cup_{i=2,3,4}\mathcal{F}_{i}])^{\top}$ (G.19)
    $=$ $\Lambda^{*}[\mathcal{B}_{1}^{*},\{1,2\}]\Phi^{*}[\{1,2\},\{1,3,4,5\}](\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\{1,3,4,5\}])^{\top}$
    It is straightforward that $\Lambda^{*}[\mathcal{B}_{1}^{*},\{1,2\}]\Phi^{*}[\{1,2\},\{1,3,4,5\}](\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\{1,3,4,5\}])^{\top}$ has rank 2 according to Condition 5 and 6. While, since $|\cup_{i=2,3,4}\mathcal{F}_{i}|=4<|\mathcal{F}_{1}|$, according to Sylvester’s rank inequality,
    $\text{rank}(\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}]\Phi[\mathcal{F}_{1},\cup_{i=2,3,4}\mathcal{F}_{i}](\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\cup_{i=2,3,4}\mathcal{F}_{i}])^{\top})$ (G.20)
    $\geq$ $\text{rank}(\Lambda[\mathcal{B}_{1}^{*},\mathcal{F}_{1}])+\text{rank}(\Phi[\mathcal{F}_{1},\cup_{i=2,3,4}\mathcal{F}_{i}])+\text{rank}(\Lambda[\cup_{i=2,3,4}\mathcal{B}_{i}^{*},\cup_{i=2,3,4}\mathcal{F}_{i}])-|\mathcal{F}_{1}|-4$
    $=$ $|\mathcal{F}_{1}|-1+4+4-|\mathcal{F}_{1}|-4$
    $=$ $3,$
    which contradicts to equation (G.19). Thus, this case does not exist.